OpenStudy (anonymous):

Please Help I'm stuck really bad can someone help me show this equation = m please: lim (((n+1)^m )-n^m) n->+inf ---------------- = m n^(m-1)

OpenStudy (anonymous):

Perhaps this helps: (1 + a)^m = 1 + m a for a<<1

OpenStudy (anonymous):

Honestly I'm just really confused by this problem, Ive tried looking at it different ways but i haven't been able to figure it out

OpenStudy (amistre64):

there is a binomial result from (a+b)^n \[\sum_{k=0}^{n}\binom{n}{k}a^{n}b^{n-k}\]

OpenStudy (amistre64):

since b=1 in this case .... that reduces it visually as \[\sum_{k=0}^{n}\binom{n}{k}a^{n}\]using n on my part was not a good idea since there is an n in your setup originally, just habit

OpenStudy (amistre64):

the binomial (n k) = n!/[(n-k)!k!] = n(n-1)(n-2)...(n-k)...(3)(2)(1) ------------------------ (n-k)...(3)(2)(1) * k! n(n-1)(n-2)...(n-k+1) ------------------- k!

OpenStudy (anonymous):

Alright so far I get it .. But Im just confused to what tipped you off to start this problem with the binomial law ?

OpenStudy (amistre64):

that (n+1)^m as m to infinity, it is a binomial

OpenStudy (amistre64):

ugh, its n to infinity lol

OpenStudy (anonymous):

Lol cool, but then the rest is still confusing though .. Are you breaking them down in parts ?

OpenStudy (amistre64):

since m is constant here, it makes the setup a little more easy to work with n^m + n^(m-1) m + n^(m-2) m(m-1)(m-2)/2! + ... - n^m --------------------------------------------------- n^(m-1)

OpenStudy (amistre64):

n^(m-1) m + n^(m-2) m(m-1)(m-2)/2! + ... ----------------------------------------- n^(m-1) m + n^(m-2-m+1) m(m-1)(m-2)/2! + n^(m-3-m+1) m(m-1)(m-2)(m-2)/3!+... m + n^(-1) m(m-1)(m-2)/2! + n^(-2) m(m-1)(m-2)(m-2)/3!+ n^(-3)... all the n parts are now fractions, and as n goes to infinty they simply zero down, and we are left with that constant first term

OpenStudy (amistre64):

\[\Large\frac{\left[\sum_{k=0}^{m}\binom{m}{k}n^k\right]-n^m}{n^{m-1}}\] \[\Large\frac{\sum_{k=0}^{m-1}\binom{m}{k}n^k}{n^{m-1}}\] \[\Large\sum_{k=0}^{m-1}\frac{\binom{m}{k}n^k}{n^{m-1}}\] \[\Large\sum_{k=0}^{m-1} \binom{m}{k}n^{k-m+1}\] looks about right to me ....

OpenStudy (anonymous):

So you literally separated the binomial part .. solved it then used it to cancel out the rest of the terms ?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

the summation notation might be seen better if n is left on the bottom: \[\Large\sum_{k=0}^{m-1} \binom{m}{k}\frac{1}{n^{m-1-k}}\] as n to infinity, all the terms zero out except for that very last term: \[\binom{m}{m-1}\frac{1}{\infty^0}:=m\]

OpenStudy (anonymous):

Alright I get it Thanks!! I just have a follow up question though .. Could we have solved the equation if we had done root of m ? I know it will cancel out all m's in the equation but I,m guessing that that's already a bad first step right ?

OpenStudy (amistre64):

i am always leary of altering the setup in such an abrupt manner, it just tends to increase the chance for error to me.

OpenStudy (amistre64):

spose we mth root it so that m is an even number, but the innards reduce to a negative; we have moved from the reals into the complex

OpenStudy (anonymous):

Yeah I agree ... Therefore it's lot simpler using the sum identities

OpenStudy (anonymous):

Thanks A lot!! You really saved me :D

OpenStudy (amistre64):

it was a fun problem to delve into :)